# ReMoS: 3D Motion-Conditioned Reaction Synthesis for Two-Person Interactions

### v2

@@ Abstract
Current approaches for 3D human motion synthesis generate high-quality animations of digital humans performing a wide variety of actions and gestures. However, a notable technological gap exists in addressing the complex dynamics of multi-human interactions within this paradigm. In this work, we present ReMoS, a denoising diffusion-based model that synthesizes full-body reactive motion of a person in a two-person interaction scenario. Assuming the motion of one person is given, we employ a combined spatio-temporal cross-attention mechanism to synthesize the reactive body and hand motion of the second person, thereby completing the interactions between the two. We demonstrate ReMoS across challenging two-person scenarios such as pair-dancing, Ninjutsu, kickboxing, and acrobatics, where one person’s movements have complex and diverse influences on the other. We also contribute the ReMoCap dataset for two-person interactions containing full-body and finger motions. We evaluate ReMoS through multiple quantitative metrics, qualitative visualizations, and a user study, and also indicate usability in interactive motion editing applications.
@@ 1 Introduction
Figure 1: Visualizations of reactive 3D motion sequences synthesized with the proposed ReMoS approach. Given the 3D motion of the acting person (red), we synthesize the 3D full-body motion of the reacting person (blue) with meaningful interactions between the two (Ninjutsu practice in the top row and Lindy Hop dancing in the middle row). The synthesized hand interactions are enlarged and highlighted with circles. Application of our results in virtual character animation (bottom row): the red arrow indicates the input action and the blue arrow indicates the synthesized reactive motions. We urge readers to watch our supplementary video.
@@
Digital 3D character motion synthesis has emerged as the next frontier for animation pipelines [49], particularly through denoising diffusion probabilistic models (DDPMs) [60]. While methods for generating character motion for various tasks such as text or music conditioned motion synthesis [18, 48, 4, 3, 34, 23, 22, 78], face and gesture synthesis [47, 71, 6, 5, 25], human-scene interaction [19, 82, 80] exist, synthesizing interactions between humans is relatively under-explored. Modeling such human-human interactions is essential for designing generative 3D human motion synthesis frameworks supporting the complex physical and social interplay of two interacting person [65]. It offers new capabilities for character animation tools and software, with applications in commercial and entertainment media [26], interactive mixed and augmented reality [14], and social robotics [74].
@@
Toward this goal, we focus on the task of modeling reactive motions. We aim to automatically generate realistic temporally aligned reactive motions of a responding person, given the continuous guiding motion of an acting individual (see Fig. 1). This alleviates a common overhead for animators, enabling them to design an acting character and automatically obtain meaningful motions for a reacting character. However, automating such a process poses two key challenges. First, two-person interaction synthesis increases the dimensionality of the already challenging problem of single-person motion synthesis. In our case, the generated reactive motions must align with the conditional signals provided in the form of 3D motion sequences from the actor. Second, doing so in a generative setting without any cues from text prompts or action labels (to obviate additional supervision and data needs) requires a careful trade-off between generating diverse motions and adhering to the narrow manifold of plausible reactions.
@@
While existing methods for synthesizing interactions [20, 10] and two-person motions [64, 42, 53] are good at generating plausible motions, they rely on additional annotations, such as action labels or text prompts, to specify the motion. Action labels typically depict actions at a high level, such as “high-five” or “salsa”, but do not capture the fine-grained synchronization with the actor’s motions. Textual descriptions offer nuance for the motions, but collecting accurate textual annotations is considerably cumbersome, particularly for contact-heavy or fast-paced two-person motions, such as dancing and exercising. Further, performing generative tasks through textual prompting remains challenging for non-AI experts [77]. Apart from the data, contact-based motion generation also needs to consider hand-based interactions to improve realism. This entails the need for frameworks capable of synthesizing hand movements with higher degrees of freedom, which are challenging to model but crucial for improving motion realism. Balancing focus between the broader full-body reactive movements and the finer hand motions involves addressing motion at two significantly different scales — a challenge exacerbated by the high degrees of freedom of the skeletons.
@@
We evaluate our approach in different scenarios, including Lindy Hop dancing, Ninjutsu, Acrobatics [24] and Kickboxing [54], and report state-of-the-art metrics (Sec. 5.3). We also perform a user study to analyze the visual quality of ReMoS compared to the ground truth and the baseline methods (Sec. 5.4).
Figure 2: Overview of ReMoS approach. Given the motion of an acting person (bottom), we automatically synthesize a plausible reactive motion for the reacting person (bottom-left). We achieve this using a denoising diffusion-based probabilistic model (centre) trained on reactive motion sequences (top-left).
@@ 2 Related Work
Synthesizing close interactions between two or multiple persons is a challenging task in animation. Early works in multi-person interaction are based on motion graphs [58], interaction patches [57], momentum-based inverse kinematics and motion blending [37], and, topologically-based pose representations [28, 29], to name a few. The increasing availability of interaction datasets [76, 54, 52, 16, 43, 39] has led to a rise in data-driven approaches for synthesizing digital partners or opponents in multi-person interaction settings. Mousas [46] uses a hidden Markov model to control the dance motions of a digital partner in an immersive setup. Ahuja et al. [1] introduce a residual-attention model to generate body pose in a conversation setting conditioned on two audio signals and the opposite person’s body pose. Starke et al. [63] propose a phase function-based network that learns asynchronous movements of each bone and its interaction with external objects. Guo et al. [24] propose the Extreme Pose Interaction dataset with a two-stream network with cross-interaction attention modules for forecasting pose sequences of two interacting characters. GAN-based models [45, 20] generate motions for an interacting person conditioned on an input character and class labels depicting the type of reaction performed. These methods leverage the daily interaction datasets such as SBU Kinect [76] and 2C [54]. InterFormer [11] uses an interaction transformer with both spatial and temporal attention to generate reactive motions given some initial seed poses of both characters. It can synthesize sparse-level interactions based on the motions of the K3HI [32] and the DuetDance [39] datasets. Concurrent to our work, Duolando [59] uses an off-policy reinforcement learning model to predict tokenized motion for a leader and follower conditioned on music. The aforementioned methods either depend on seed motion as input or require additional conditions to drive the motion. In contrast, ReMoS focuses on synthesizing well-synchronized full-body and hand motions for the reactor conditioned only on the actor’s 3D motions, without using any additional labels or prompts.
@@
Denoising diffusion models [60] have recently demonstrated their high potential in generative human motion modeling, specifically in single-person conditional motion synthesis. Conditional single-person motion generation has been performed using diffusion-based approaches for co-speech gestures [2, 84], audio-driven motion [67, 13, 83], and text-driven motion [79, 66, 53, 75] synthesis. Diffusion-based techniques have also been extensively used for conditional synthesis for human-object interactions [40, 72, 41, 38], hand-object interactions [73, 44], and human-scene interactions [33]. Guided motion diffusion models such as GMD [35] and TraceAndPace [51] incorporate spatial constraints on motion trajectories, to guide the motion towards a goal at inference time. OmniControl [70] proposes spatial and realism guidance to control any joint for diffusion-based human motion generation. For two-person interaction synthesis, RAIG [64] proposes a diffusion-based, role-aware approach for two-person interactions, given separate textual descriptions for each person. BiGraphDiff [10] uses graph transformer denoising diffusion to learn two-person interaction conditioned on action labels. ComMDM [53] uses a communication block between two MDMs [66] to coordinate two-person interaction generation. ContactGen [21] presents a diffusion-based contact prediction module that adaptively estimates potential contact regions between two humans according to the interaction label. Liang et al. [42] proposes InterGen, a diffusion-based model to synthesize two-person interactions given text descriptions as input conditions. They also propose the InterHuman dataset that includes diverse two-person interaction scenarios with rich text annotations. Inspired by these recent approaches, we design a DDPM-based approach for 3D motion-conditioned reaction synthesis. Our task is to implicitly learn the fine-grained synchronization that is possible between two interacting persons, just from one of the person’s motions, and synthesize the corresponding movements of the second person. To do so, we propose ReMoS, a cascaded diffusion model with a combined spatio-temporal cross-attention mechanism that helps in learning the fine-grained inter-person dependencies across time, without the requirement for additional prompts or labels, thus going beyond the state-of-the-art methods [42, 64] for two-person motion synthesis. ReMoS further synthesizes plausible hand motions for the reactor to incorporate realistic hand-based interactions.
@@ 3.1 DDPM for Reactive Motion Synthesis
$X^{\left(t\right)}=\sqrt{\bar{\alpha}_{t}}X^{\left(0\right)}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},\;\;\epsilon\sim\mathcal{N}\left(0,I\right),$ (1)
where $\bar{\alpha}_{t}$ controls the rate of diffusion and $t\in[0,T]$. Reversing this diffusion process allows for sampling novel motion sequences from a multivariate Gaussian distribution $\mathit{p}\left(X^{\left(T\right)}\right)\sim\mathcal{N}\left(0,I\right)$ as
@@
Following standard practice, we approximate the computationally intractable term $\mathit{p}\left(X^{\left(t-1\right)}|X^{\left(t\right)}\right)$ with a learnable function $\mathit{f_{\theta}}\left(X^{\left(t\right)},t\right)$, and optimize $\theta$ to encode the space of human reactive motions. In our case, the reactive motion $X$ is also conditioned on the actor’s motion $Y$. Therefore, we modify the learnable function as $X^{(t-1)}=\mathit{f}_{\theta}\left(X^{\left(t\right)},t,Y\right)$. Following recent works [50, 66], we directly estimate the original motion $X^{(0)}=\mathit{f}_{\theta}\left(X^{\left(t\right)},t,Y\right)$ from our diffusion model.
@@ 3.2 ReMoS Framework
$X_{B}^{(0)}$ $=\mathit{f}_{{\theta}_{B}}\left(X_{B}^{(t)},t,Y_{B}\right),$ (3)
$X_{H}^{(0)}$ $=\mathit{f}_{{\theta}_{H}}\left(X_{H}^{(t)},t,Y_{H},\mathbbm{1}_{H}\left(X_{B}^{(0)},Y_{B}\right)\right),\text{and}$ (4)
$X^{(0)}$ =$=\{X_{B}^{(0)},X_{H}^{(0)}\},$ (5)
where $\mathbbm{1}_{H}$ is a binary, spatio-temporal hand-interaction mask function that determines which hand joints of the reactor interacts with the actor at any given frame. Our cascaded framework comes from the observation that full-body articulations occur at significantly different scales than hand articulations. As the body motions affect the hand articulations, our cascaded framework feeds the generated body motions as an additional condition to the hand generation module (Fig. 3). We also train the body and the hand generation modules separately, with additional hand-interaction-aware attention mapping for the hand joints. Previous work [19] used such disjoint training strategy on conditional motion synthesis for single persons. To better accommodate for two-person interactions in our case, we benefit from using a cascaded diffusion strategy [31].
Figure 3: ReMoS Framework. Given the full-body sequence of the actor (in red, left), we input noisy body and hand samples (from below) in a cascaded fashion. We synthesize the body samples first, and use them for hand-interaction-aware attention masking (top-center) to synthesize the denoised hand samples (top-right). The full-body reactive motion is a concatenation of the denoised body and hand samples (in blue, right).
@@
In the first stage of our cascaded framework, we diffuse the reactor’s body motion, $X_{B}^{(t)}$, at each diffusion step $t$, and feed it into a transformer decoder [68] block. To condition the reactor’s body motion on the actor’s body motion $Y_{B}$, we introduce a combined spatio-temporal cross-attention (CoST-XA) mapping. CoST-XA combines the spatial and temporal features of the actor’s and the reactor’s motions through an attention matrix to simultaneously learn the inter-dependencies between each pair of reactor and actor body joints $\left(x_{n,j},y_{n,j}\right)$, $x_{n,j}\in X_{B}^{(t)}$ and $y_{n,j}\in Y_{B}$, respectively, at each frame. We denote the query from the reactor’s motion features as $Q_{B}\in\mathbb{R}^{(NJ_{B})\times 3}$ and the key, value pair from the actor’s motion features as $K_{B}\in\mathbb{R}^{(NJ_{B})\times 3}$ and $V_{B}\in\mathbb{R}^{(NJ_{B})\times 3}$, respectively. The formulation above leads to an attention matrix of dimensions $(NJ_{B}){\times}(NJ_{B})$. This is in contrast to previous approaches [64, 42], where the methods use cross-attention only on the temporal features of the motion leading to an attention matrix of shape $N\times N$. InterFormer [11] sequentially uses spatial and temporal cross-attention modules, instead of combining them; this may result in a partial loss of information on fine-grained, inter-person interactions across time, especially in the absence of additional annotations (prompts or action labels) to condition the reactions. In contrast, our proposed cross-attention module, which we define as
@@
crucially considers combinations of spatial and temporal interaction features between different body segments of the actor and the reactor to efficiently synchronize their motions. We simultaneously project the diffusion timestep $t$ at each denoising step into the transformer block after the attention blocks and use a final layer of fully-connected network with SiLU activations [15] and batch normalization [7] to generate the body motions. The first stage output is a vector $X_{B}^{(0)}\in\mathbb{R}^{N\times J_{B}\times 3}$ representing the reactor’s clean, synthesized body motion.
@@
The second stage of our cascaded framework synthesizes the reactor’s hand joints. We input noisy samples $X_{H}^{(t)}$ into a similar transformer decoder block that is conditioned on the actor’s hand motions $Y_{H}$. While generating the reactive hand motions, we require the attention weights to be high for the hands used in the hand-based interactions, and low for the passive hands. We explicitly enforce this using a spatio-temporal hand-interaction-aware cross-attention mapping (H-XA) to encourage learning localized interaction features. For H-XA to work, we introduce the binary hand-interaction mask $\mathbbm{1}_{H}\in\left\{0,1\right\}^{N\times 4}$ with its entries calculated by thresholding the distances of the actor’s and reactor’s wrist joints in $Y_{B}$ and $X_{B}^{(0)}$, respectively to the nearest body joints of the other person. The active entries of $\mathbbm{1}_{H}$ determine which of the actor’s and reactor’s hands are sufficiently close to the body of the other person (and hence interacting). We then project the encoded $Y_{H}$ features into each decoder layer using H-XA as
H-XA $=\operatorname{softmax}\left(\mathbbm{1}_{H}\cdot\frac{Q_{H}K_{H}^{T}}{\sqrt{d_{K_{H}}}}\right)V_{H},$ (7)
where $K_{H}$ and $V_{H}$ are the hand motion features from the actor, and $Q_{H}$ are the hand motion features from the reactor, with $d_{K_{H}}$ channels. We project the diffusion timestep $t$ at each denoising step into the transformer block and use fully-connected networks with SiLU activations to generate the clean, synthesized reactor hand motions $X_{H}^{(0)}\in\mathbb{R}^{N\times J_{H}\times 3}$. At the end, to obtain the full reactive motion $X^{(0)}$, we concatenate $\{X_{B}^{(0)},X_{H}^{(0)}\}$ as in Eqn. (5).

### v3

@@ Abstract
Current approaches for 3D human motion synthesis generate high-quality animations of digital humans performing a wide variety of actions and gestures. However, a notable technological gap exists in addressing the complex dynamics of multi-human interactions within this paradigm. In this work, we present ReMoS, a denoising diffusion-based model that synthesizes full-body reactive motion of a person in a two-person interaction scenario. Given the motion of one person, we employ a combined spatio-temporal cross-attention mechanism to synthesize the reactive body and hand motion of the second person, thereby completing the interactions between the two. We demonstrate ReMoS across challenging two-person scenarios such as pair-dancing, Ninjutsu, kickboxing, and acrobatics, where one person’s movements have complex and diverse influences on the other. We also contribute the ReMoCap dataset for two-person interactions containing full-body and finger motions. We evaluate ReMoS through multiple quantitative metrics, qualitative visualizations, and a user study, and also indicate usability in interactive motion editing applications. More details are available on the project page: https://vcai.mpi-inf.mpg.de/projects/remos.
Keywords:
Reactive Motion Synthesis Denoising Diffusion Model
@@ 1 Introduction
Digital 3D character motion synthesis has emerged as the next frontier for animation pipelines [50], particularly through denoising diffusion probabilistic models (DDPMs) [61]. While methods for generating character motion for various tasks such as text or music conditioned motion synthesis [18, 49, 4, 3, 34, 23, 22, 80], face and gesture synthesis [48, 73, 6, 5, 25, 47], human-scene interaction [19, 84, 82, 71] exist, synthesizing interactions between humans is relatively under-explored. Modeling such human-human interactions is essential for designing generative 3D human motion synthesis frameworks supporting the complex physical and social interplay of two interacting persons [66]. It offers new capabilities for character animation tools and software, with applications in commercial and entertainment media [26], interactive mixed and augmented reality [14], and social robotics [76].
Figure 1: Visualizations of reactive 3D motion sequences synthesized with the proposed ReMoS approach. We synthesize the 3D full-body motion of the reactor (blue) conditioned only on the 3D motion of the actor (red), thereby completing the interactions between the two (Ninjutsu practice on the left and Lindy Hop dancing on the right). The synthesized hand interactions are enlarged and highlighted with circles.
@@
Toward this goal, we focus on the task of modeling reactive motions. We aim to automatically generate realistic, temporally-aligned reactive motions of a responding person given the continuous motion of a guiding person (see Fig. 1). This alleviates a common overhead for animators, enabling them to design an acting character and automatically obtain meaningful motions for a reacting character. However, automating such a process poses two key challenges. First, two-person interaction synthesis increases the dimensionality of the already challenging problem of single-person motion synthesis. In our case, the generated reactive motions must align with the conditional signals provided in the form of 3D motion sequences from the actor. Second, doing so in a generative setting without any cues from text prompts or action labels (to obviate additional supervision and data needs) requires a careful trade-off between generating diverse motions and adhering to the narrow manifold of plausible reactions.
@@
While existing methods for synthesizing interactions [20, 10] and two-person motions [65, 42, 54] are good at generating plausible motions, they rely on additional annotations, such as action labels or text prompts, to specify the motion. Action labels typically depict actions at a high level, such as “high-five” or “salsa”, but do not capture the fine-grained synchronization with the actor’s motions. Textual descriptions offer nuance for the motions, but collecting accurate textual annotations is considerably cumbersome, particularly for contact-heavy or fast-paced two-person motions, such as dancing and exercising. Further, performing generative tasks through textual prompting remains challenging for non-experts [79]. Apart from the data, contact-based motion generation also needs to consider hand-based interactions to improve realism. This entails the need for frameworks capable of synthesizing hand movements with higher degrees of freedom, which are challenging to model but crucial for improving motion realism. Balancing focus between the broader full-body reactive movements and the finer hand motions involves addressing motion at two significantly different scales — a challenge exacerbated by the high degrees of freedom of the skeletons.
@@
We evaluate our approach in different scenarios, including Lindy Hop dancing, Ninjutsu, Acrobatics [24] and Kickboxing [55], and report state-of-the-art performance (Sec. 5.3). We also report a user study on comparing the visual quality of ReMoS compared to the ground truth and the baseline methods (Sec. 5.4).
@@ 2 Related Work
Synthesizing close interactions between two or multiple persons is a challenging task in animation. Early works in multi-person interaction are based on motion graphs [59], interaction patches [58], momentum-based inverse kinematics and motion blending [37], and topologically-based pose representations [28, 29], to name a few. The increasing availability of interaction datasets [78, 55, 53, 16, 43, 39] has led to a rise in data-driven approaches for synthesizing digital partners or opponents in multi-person interaction settings. Mousas [46] uses a hidden Markov model to control the dance motions of a digital partner in an immersive setup. Ahuja et al. [1] introduce a residual-attention model to generate body pose in a conversation setting conditioned on two audio signals and the opposite person’s body pose. Starke et al. [64] propose a phase function-based network that learns asynchronous movements of each bone and its interaction with external objects. Guo et al. [24] propose the Extreme Pose Interaction dataset with a two-stream network with cross-interaction attention modules for forecasting pose sequences of two interacting characters. GAN-based models [45, 20] generate motions for an interacting person conditioned on an input character and class labels depicting the type of reaction performed. These methods leverage the daily interaction datasets such as SBU Kinect [78] and 2C [55]. InterFormer [11] uses an interaction transformer with both spatial and temporal attention to generate reactive motions given some initial seed poses of both characters. It can synthesize sparse-level interactions based on the motions of the K3HI [32] and the DuetDance [39] datasets. Concurrent to our work, Duolando [60] uses an off-policy reinforcement learning model to predict tokenized motion for a leader and follower conditioned on music. The aforementioned methods either depend on seed motion as input or require additional conditions to drive the motion. In contrast, ReMoS focuses on synthesizing well-synchronized full-body and hand motions for the reactor conditioned only on the actor’s 3D motions, without using any additional labels or prompts.
@@
Denoising diffusion models [61] have recently demonstrated their high potential in generative human motion modeling, specifically in single-person conditional motion synthesis. Conditional single-person motion generation has been performed using diffusion-based approaches for co-speech gestures [2, 86], audio-driven motion [68, 13, 85], and text-driven motion [81, 67, 54, 77] synthesis. Diffusion-based techniques have also been extensively used for conditional synthesis for human-object interactions [40, 74, 41, 38], hand-object interactions [75, 44], and human-scene interactions [33]. Guided motion diffusion models such as GMD [35] and TraceAndPace [52] incorporate spatial constraints on motion trajectories, to guide the motion towards a goal at inference time. OmniControl [72] proposes spatial and realism guidance to control any joint for diffusion-based human motion generation. For two-person interaction synthesis, RAIG [65] proposes a diffusion-based, role-aware approach for two-person interactions, given separate textual descriptions for each person. BiGraphDiff [10] uses graph transformer denoising diffusion to learn two-person interaction conditioned on action labels. ComMDM [54] uses a communication block between two MDMs [67] to coordinate two-person interaction generation. ContactGen [21] presents a diffusion-based contact prediction module that adaptively estimates potential contact regions between two humans according to the interaction label. Liang et al. [42] propose InterGen, a diffusion-based model to synthesize two-person interactions given text descriptions as input conditions. They also propose the InterHuman dataset that includes diverse two-person interaction scenarios with rich text annotations. Inspired by these recent approaches, we design our proposed method ReMoS as a DDPM-based 3D motion-conditioned reaction synthesis method, consisting of a cascaded diffusion model with a combined spatio-temporal cross-attention mechanism. This allows ReMoS to implicitly learn the fine-grained synchronization between two interacting persons from only one of the person’s motions, without any need for additional prompts or labels, and synthesize the corresponding motions of the second person. ReMoS further synthesizes plausible hand motions for the reactor to incorporate realistic hand-based interactions.
Figure 2: ReMoS Overview. Given the motion of the actor (bottom-middle, in red), we synthesize a plausible motion for the reactor (bottom-left, in blue). We achieve this using a denoising diffusion-based probabilistic model (center) trained on reactive motion sequences (top-left, in blue).
@@ 3.1 DDPM for Reactive Motion Synthesis
$X^{\left(t\right)}=\sqrt{\bar{\alpha}_{t}}X^{\left(0\right)}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},\qquad\epsilon\sim\mathcal{N}\left(0,I\right),$ (1)
where $\bar{\alpha}_{t}$ controls the rate of diffusion and $t\in\left[0,T\right]$. Reversing this diffusion process allows for sampling novel motion sequences from a multivariate Gaussian distribution $\mathit{p}\left(X^{\left(T\right)}\right)\sim\mathcal{N}\left(0,I\right)$ as
@@
We approximate the computationally intractable term $\mathit{p}\left(X^{\left(t-1\right)}|X^{\left(t\right)}\right)$ with a learnable function $\mathit{f_{\theta}}\left(X^{\left(t\right)},t\right)$, and optimize $\theta$ to encode the space of human reactive motions. Our reactive motion $X$ is also conditioned on the actor’s motion $Y$. Therefore, we modify the learnable function as $X^{\left(t-1\right)}=\mathit{f}_{\theta}\left(X^{\left(t\right)},t,Y\right)$. Following recent works [51, 67], we estimate the original motion $X^{\left(0\right)}=\mathit{f}_{\theta}\left(X^{\left(t\right)},t,Y\right)$ from our diffusion model by iterating through all $t$ denoising steps during inference.
@@ 3.2 ReMoS Framework
$X_{B}^{\left(0\right)}$ $=\mathit{f}_{{\theta}_{B}}\left(X_{B}^{\left(t\right)},t,Y_{B}\right),$ (3)
$X_{H}^{\left(0\right)}$ $=\mathit{f}_{{\theta}_{H}}\left(X_{H}^{\left(t\right)},t,Y_{H},\mathbbm{1}_{H_{A}}\left(Y_{B}\right)\mathbbm{1}_{H_{R}}\left(X_{B}^{\left(0\right)}\right)\right),\quad\text{and}$ (4)
$X^{\left(0\right)}$ =$=\left\{X_{B}^{\left(0\right)},X_{H}^{\left(0\right)}\right\},$ (5)
where $\mathbbm{1}_{H_{A}}$ and $\mathbbm{1}_{H_{R}}$ are binary, spatio-temporal hand-interaction mask functions that determine which hand joints of the reactor and actor interact at any given frame. Our cascaded framework comes from the observation that full-body articulations occur at significantly different scales than hand articulations. Since the body motions affect the hand articulations, our cascaded framework feeds the generated body motions as an additional condition to the hand generation module (Fig. 3). We also train the body and the hand generation modules separately, with additional hand-interaction-aware attention mapping for the hand joints. Previous work [19] used such disjoint training strategy on conditional motion synthesis for single persons. To better accommodate for two-person interactions in our case, we benefit from using a cascaded diffusion strategy [31].
Figure 3: ReMoS Framework. Given the full-body sequence of the actor (left, in red), we input noisy body and hand samples (from below) in a cascaded fashion. We synthesize the body samples first, and use them for hand-interaction-aware attention masking (top-center) to synthesize the denoised hand samples (top-right). The full-body reactive motion is a concatenation of the denoised body and hand samples (right, in blue).
@@
In the first stage of our cascaded framework, we diffuse the reactor’s body motion, $X_{B}^{\left(t\right)}$, at each diffusion step $t$, and feed it into a transformer decoder [69] block. To condition the reactor’s body motion on the actor’s body motion $Y_{B}$, we introduce a combined spatio-temporal cross-attention (CoST-XA) mapping. CoST-XA combines the spatial and temporal features of the actor’s and the reactor’s motions through an attention matrix to simultaneously learn the inter-dependencies between each pair of reactor and actor body joints $\left(x_{n,j},y_{n,j}\right)$, $x_{n,j}\in X_{B}^{\left(t\right)}$ and $y_{n,j}\in Y_{B}$, respectively, at each frame. Denoting the query from the reactor’s motion features as $Q_{B}\in\mathbb{R}^{NJ_{B}\times 3}$ and the key-value pair from the actor’s motion features as $K_{B}\in\mathbb{R}^{NJ_{B}\times 3}$ and $V_{B}\in\mathbb{R}^{NJ_{B}\times 3}$, respectively, we formulate an attention matrix of dimensions $NJ_{B}\times NJ_{B}$. In contrast, most previous approaches [65, 42] use cross-attention only on the temporal features of the motion using an attention matrix of shape $N\times N$. InterFormer [11] sequentially uses spatial and temporal cross-attention modules instead of combining them, which may result in a partial loss of information on fine-grained, inter-person interactions across time, especially in the absence of additional annotations (prompts or action labels) to condition the reactions. Our proposed cross-attention module, which we define as
@@
crucially considers combinations of spatial and temporal interaction features between different body segments of the actor and the reactor to efficiently synchronize their motions. We simultaneously project the diffusion timestep $t$ at each denoising step into the transformer block after the attention blocks and use a final layer of fully-connected network with SiLU activations [15] and batch normalization [7] to generate the body motions. The first stage output is a vector $X_{B}^{\left(0\right)}\in\mathbb{R}^{N\times J_{B}\times 3}$ representing the reactor’s clean, synthesized body motion.
@@
The second stage of our cascaded framework synthesizes the reactor’s hand joints. We input noisy samples $X_{H}^{\left(t\right)}$ into a similar transformer decoder block conditioned on the actor’s hand motions $Y_{H}$. Here, we require the attention weights to be high for the hands used in the hand-based interactions and low for the passive hands. We explicitly enforce this using a spatio-temporal hand-interaction-aware cross-attention mapping (H-XA) to encourage learning localized interaction features. For H-XA to work, we introduce binary hand-interaction masks $\mathbbm{1}_{H_{A}}$ and $\mathbbm{1}_{H_{R}}\in\mathbb{R}^{N\times J_{H}}$. Its entries are calculated by thresholding the distances of the actor’s and reactor’s wrist joints in $Y_{B}$ and $X_{B}^{\left(0\right)}$, respectively, to the nearest body joints of the other person. The active entries of $\mathbbm{1}_{H_{A}}$ and $\mathbbm{1}_{H_{R}}$ determine which of the actor’s and reactor’s hands are sufficiently close to the body of the other person (and hence interacting). We then project the encoded $Y_{H}$ features into each decoder layer using H-XA, as
[…]
where $K_{H}$ and $V_{H}$ are the hand motion features from the actor, $Q_{H}$ are the hand motion features from the reactor with $d_{K_{H}}$ channels, and $\odot$ denotes element-wise product for masking the query and key values. We project the diffusion timestep $t$ at each denoising step into the transformer block and use fully-connected networks with SiLU activations to generate the clean, synthesized reactor hand motions $X_{H}^{\left(0\right)}\in\mathbb{R}^{N\times J_{H}\times 3}$. At the end, to obtain the full reactive motion $X^{\left(0\right)}$, we concatenate $\left\{X_{B}^{\left(0\right)},X_{H}^{\left(0\right)}\right\}$ as in Eqn. 5.